Victor Co. sells ratios for $30 each. Fixed costs total $15,000. Unit variable costs $20. The tax rate is 40%. How many radios must Victor Co. sell to earn after-tax profit of $21,000?
Answer #1

Profit before tax = 21000/(1-0.40) = $35000

Units need to sale = (Fixed cost+ Profit)/(sale - variable cost)

= (15000+35000)/(30-20)

= 5000 units